OpenStudy (anonymous):

solve this ∫ 1/(√(5x+2) + √(2x+1) ) dx

OpenStudy (anonymous):

No, it's all under the denominator, mihirb.

OpenStudy (anonymous):

oh it is didnt notice that

OpenStudy (anonymous):

yes its all in denominator

OpenStudy (anonymous):

$$\int \frac{1}{\sqrt {5x + 2} + \sqrt {2x + 1}} \, dx$$

OpenStudy (anonymous):

oh srry

OpenStudy (anonymous):

yeah thats the question

OpenStudy (anonymous):

It's a little bit weird, but u-substitution works.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

i tried but i m not getting it..

OpenStudy (anonymous):

First, rationalize the integrand by multiplying by the conjugate of the denominator. Can you do that?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i m getting after rationalizing\[\frac{ \sqrt{5x+2} +\sqrt{2x+1} }{ 3x+3 }\]

OpenStudy (anonymous):

sorry there should be -ve sign in numerator

OpenStudy (anonymous):

Yes, now use partial fractions to split up the integral.

OpenStudy (anonymous):

done.. then?

OpenStudy (anonymous):

and 3x+1 in the denominator

OpenStudy (anonymous):

No, it's -3x -1.

OpenStudy (anonymous):

what i did't understand?

OpenStudy (anonymous):

multiplying by conjugate in the denominator should lead to 5x+2-(2x+1) which is 3x+1 if you distribute the negative properly

OpenStudy (anonymous):

Which is just the same thing as a minus sign over the whole thing, which is a minus sign in the denominator. Er, well, this is what I'm using, so go with this for clarity, it's just a sign mix-up: $$\int \frac {\sqrt {2x + 1}} {-3x - 1}\, dx - \int \frac {\sqrt {5x + 2}} {-3x - 1}\, dx$$

OpenStudy (anonymous):

what after this\[\frac{ \sqrt{5x+2} }{ 3x+3 } - \frac{ \sqrt{2x+1} }{ 3x+3 }\]

OpenStudy (anonymous):

okay then?

OpenStudy (anonymous):

Substitute \(u = 5x + 2\) and its corresponding \(du = \frac{5}{2\sqrt {5x + 2}}\).

OpenStudy (anonymous):

We're starting with that integral because it's a bit easier

OpenStudy (anonymous):

if u=5x+2 then how u got the value of du?

OpenStudy (anonymous):

Sorry, I meant \(u = \sqrt {5x + 2}; du = \frac {5}{2\sqrt{5x + 2}} dx\).

OpenStudy (anonymous):

okay .. now i got it

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

We have: $$-\frac{2}{5} \int \frac{u^2}{-\frac 3 5 (u^2 - 2) - 1} \, du$$ Now, just cancel out stuff with a bit of interesting manipulation. $$-\frac{2}{5} \int -\frac{5u^2}{3u^2 - 1} \, du$$

OpenStudy (anonymous):

Now factor out all of those constants. Move them to other integral if you please.

OpenStudy (anonymous):

What should i do now?? and what abt the other term?

OpenStudy (anonymous):

@dan815 u got it wrong.. the whole expression is in denominator..

OpenStudy (anonymous):

see it in one the comments..

OpenStudy (dan815):

oh ok

OpenStudy (dan815):

like this?

OpenStudy (dan815):